OpenStudy (anonymous):

\frac{ 2 }{ 3a} \times \frac{ 2 }{ a }

OpenStudy (anonymous):

@twilightrain

OpenStudy (anonymous):

\frac{ 2 }{ 3a} \times \frac{ 2 }{ a }

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

\frac{ 2 }{ 3a} \times \frac{ 2 }{ a^2 }?

OpenStudy (anonymous):

make a new question like i said